R2. The goal in this problem is to create something called a P-P plot (youll need the result from R1 at a key place). Heres the idea: We start by choosing (say) 100 values for q evenly spaced between 0 and 10 (qvals). We also generate (say) 100000 random numbers from what R claims is an exponential distribution with (say) λ-1/2. Using a for loop, for each value of q, we calculate what percentage of numbers from the exponential distribution we expect to be less than q (see R1), and also, what percentage of numbers from our list of 100000 that actually are less than q. After doing all this, we plot the 100 expected percentages against the 100 actual percentages Think about what graph we should get before you complete the code and then run the below code. Include a rough sketch of your plot and explain why it is shaped the way it is. 1 set.seed (3.1415) # bring our random nesses into alignment # how many values for q 3 ngs 4 lambda = 1/2 6 expectedpercentages - 7 actualpercentages 8 nnums - 9 nums # use seq function to help you # set up vector for expected values # set up vector for actual values # how many random numbers to create # make the random nums 10 11 for i in 1:nqs) 12 expectedpercentages[i]- 13 actualpercentagesi] 14 15 16 plot
Answer #1

If the actual percentages match the theoretical percentages , the points in a PP Plot plot would fall on a 45 line.

Here we have drawn a sample from what R claims to be an exponential distribution, and comparing it against actual exponential distribution. Hence the actual percentages should closely match with that of the theoretical/expected percentages and the points would be around 45 line.

If x has exponential distribution with parameter lambda=rac{1}{2} , the cdf of X is

P(X<x)=F(x)=1-e^{lambda x}=1-e^{-rac{x}{2}}

The theoretical percentage of numbers from exponential distribution we expect to be less than q is

Expected Percentages = P(X < g) x 100 = (1-e-2) × 100

Following is the R code (all statements starting with # are comments)

#set the random seed set.seed (3.1415) #set number of values for q nqs<-100 # 3et the parameter for exponential di 3 tribution lambda<-1/2 #gene rate nqs values of q equally spaced between 0 and 10 qva13(-seq (from=0, to-10, length. out=nq3 ) #set up the vector of expected values expectedpercentages<-numeric (nqs) #set up the vector of actual values actualpercentages<-numeric (nqS) #set the number of random numbers to create nnurns<-100000 #generate nnums random numbers from exponential distribution nums<-rexp (nnums, lambda) #find the percentages for each value in qvals for (i in 1:nqs) i #find the pr (Kqva13[1]) * 100 expectedpercentages [1] <-(1-exp (-qvals [1] /2) ) *100 #actual percentage, count the number from exponential actualpercentages [1](-sum (nums<=qvals [1] ) /nnums * 100 distribution less than qva13[1] plot (actualpercentages, expectedpercentages, xlab= Actual abline (0,1, col=red) Percentages, ylab=Expected Percentages,main=P-P plott

output is

P-P plot 100 80 40 60 20 Actual Percentages

We can say that the numbers generated from what R claims to be exponential, are indeed exponential as the points lie on a 45 line